### v2

# Infinitely many non-collapsed steady Ricci solitons on complex line bundles

Abstract.
We construct a continuous 3-parameter family of non-shrinking Ricci solitons on complex line bundles $\mathcal{O}(k)$ over $\mathbb{CP}^{2m+1}$, where the base space is not necessarily Kähler–Einstein. Each $\mathcal{O}(k)$ with $k\in[3,2m+1]$ admits at least one asymptotically conical (AC) Ricci-flat metric in this family. For each $\mathcal{O}(k)$ with $k\geq 3$, the family includes infinitely many asymptotically paraboloidal (AP) steady Ricci solitons.
2020 Mathematics Subject Classification: 53E20 (primary), 53C25 (secondary). Keywords: Ricci soliton, cohomogeneity one metric. The author is supported by the NSFC (No. 12301078), the Natural Science Foundation of Jiangsu Province (BK-20220282), and XJTLU Research Development Fund (RDF-21-02-083).
1. Introduction
Ricci solitons are candidates for blow-up singularities of the Ricci flow. Specifically, a shrinking Ricci soliton is a candidate for a Type I singularity while a steady Ricci soliton is a candidate for a Type II singularity. From Perelman’s no-collapsing theorem [Per02], Type II singularities must be non-collapsed.
Complex line bundles $\mathcal{O}(k)$ over $\mathbb{CP}^{n}$ provide a rich source of examples for gradient Ricci solitons. The principal orbit of $\mathcal{O}(k)$ is the lens space $\mathbb{S}^{2n+1}/\mathbb{Z}_{k}$, a circle bundle over $\mathbb{CP}^{n}$. The integer $k$ classifies the circle bundle, corresponding to a multiple of the indivisible element in $H^{2}(\mathbb{CP}^{n},\mathbb{Z})$. This setup has led to numerous examples of Ricci solitons with various asymptotic behaviors. The Bérard-Bergery metric on $\mathcal{O}(n+1)$ is an AC Ricci-flat metric; see [BB82]. By [Cao96], it is known that $\mathcal{O}(n+1)$ admits a 1-parameter family of steady Kähler–Ricci solitons. These solitons have collapsed volume growth. In particular, they are asymptotically cigar-paraboloidal (ACP). Expanding Kähler–Ricci solitons on $\mathcal{O}(k)$ with $k>n+1$ are found in [FIK03]. The $U(n+1)$-invariant ansatz is further generalized to complex line bundles over products of Fano manifolds in [DW11]. Recently, new steady Ricci solitons were found on $\mathcal{O}(k)$ independently by [Win23], [Sto24] and [App17]. It was proved in [App17] that each $\mathcal{O}(k)$ with $k>n+1$ admits a non-collapsed AP steady Ricci soliton.
In this article, we examine complex line bundles $\mathcal{O}(k)$ over $\mathbb{CP}^{2m+1}$, where the symmetry condition is relaxed from $U(2m+2)$-invariance to $Sp(m+1)Sp(1)$-invariance, allowing the base space to be squashed. Specifically, the associated group triple is:
(1.1) $(\mathsf{K},\mathsf{L},\mathsf{G})=([Sp(m)\Delta U(1)]\times\mathbb{Z}_{k},Sp(m)U(1)U(1),Sp(m+1)U(1)).$
Identify $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$ as the quotient $\{\mathbf{q}\mid\mathbf{q}\in\mathbb{H}^{m+1},|\mathbf{q}|=1\}/\{\mathbf{q}\sim e^{\frac{2\pi i}{k}}\mathbf{q}\}$. Each element $(A,z)\in Sp(m+1)U(1)$ acts on $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$ via $\mathbf{q}\mapsto A\mathbf{q}\bar{z}$. The isotropy subgroup of $(0,\dots,0,1)$ is
$\mathsf{K}=\left\{\left(\mathrm{diag}(A,e^{\frac{2l\pi i}{k}}z),z\right)\mid z\in U(1),l=0,...,k-1\right\}.$
The group action descends to $\mathbb{CP}^{2m+1}$, where the isotropy subgroup of [0::0:1][0:...:0:1] is
$\mathsf{H}=\left\{\left(\mathrm{diag}(A,z_{1}),z_{2}\right)\mid z_{1},z_{2}\in U(1)\right\}.$
In the special case where $k=1$, this construction yields the quaternionic Hopf fibration.
Since the principal orbit is locally isometric to $\mathbb{S}^{4m+3}$, the cohomogeneity one dynamical system on $\mathcal{O}(k)$ is the same for different values of $k$. The same dynamical system also appeared in the author’s earlier work [Chi26], where a continuous 1-parameter family of non-collapsed steady Ricci solitons were constructed on Euclidean spaces. However, the geometric interpretation of the solutions is different in the present setting. Here we study complex line bundles $\mathcal{O}(k)\to\mathbb{CP}^{2m+1}$, where the global existence of complete metrics depends crucially on the integer parameter $k$. This leads to new families of complete Ricci solitons.
For $k\in\{1,2\}$, we construct a compact invariant set to obtain the following theorem.
Theorem 1.1.
For $k\in\{1,2\}$, there exists a continuous $3$-parameter family of complete Ricci solitons $\{\xi(k,\theta,s_{4},s_{5})\mid\theta\in(0,\pi),s_{4}\geq 0,s_{5}\geq 0\}$ on $\mathcal{O}(k)$. In particular,
  1. (1)
    each $\xi(k,\theta,s_{4},s_{5})$ with $s_{4},s_{5}>0$ is an AC non-Einstein expanding Ricci soliton;
  2. (2)
    each $\xi(k,\theta,0,s_{5})$ with $s_{5}>0$ is an asymptotically hyperbolic (AH) Einstein metric;
  3. (3)
    each $\xi(k,\theta,0,0)$ is an asymptotically locally conical (ALC) Ricci-flat metric. The base of the asymptotic cone is the non-Kähler $\mathbb{CP}^{2m+1}/\mathbb{Z}_{k}$;
  4. (4)
    each $\xi(k,\theta,s_{4},0)$ with $s_{4}>0$ is an ACP steady Ricci soliton. The base of the asymptotic paraboloid is the non-Kähler $\mathbb{CP}^{2m+1}/\mathbb{Z}_{k}$.
The parameter $s_{5}$ controls the generalized mean curvature of the principal orbit. As the generalized mean curvature is freely scaled under homothetic change if $\epsilon=0$, the parameter $s_{5}$ is not needed and we set $s_{5}=0$ for the steady case. The parameter $s_{4}$ is related to the second derivative of the potential function, and the Ricci soliton is Einstein if $s_{4}=0$. The parameter $\theta$ controls how $\mathbb{CP}^{m+1}$ is squashed as a twistor space over $\mathbb{HP}^{m}$. Solutions with $\theta=0$ admit $U(2m+2)$-symmetry and we have the Fubini–Study $\mathbb{CP}^{m+1}$. The limiting case where $\theta=\pi$ blows up the base space $\mathbb{HP}^{m}$. The cohomogeneity one steady soliton equation of this limiting case is equivalent to that of a complex line bundle over an irreducible $\mathbb{CP}^{1}$, whose principal orbit is $\mathbb{S}^{3}/\mathbb{Z}_{k}$.
If $k\geq 3$, our techniques for proving Theorem 1.1 become obsolete. Nevertheless, we generalize [App17, Theorem 5.1] and obtain the following existence theorems.
Theorem 1.2.
For each $k\geq 3$ and $\theta\in(0,\pi)$, there exists a $\beta_{k,\theta}\geq 0$ such that $\{\xi(k,\theta,s_{4},s_{5})\mid s_{4}\geq\beta_{k,\theta},s_{5}>0\}$ is a continuous $2$-parameter family of complete expanding Ricci solitons. In particular,
  1. (1)
    each $\xi(k,\theta,s_{4},s_{5})$ with $s_{4}>\beta_{k,\theta}$ is an AC non-Einstein Ricci soliton;
  2. (2)
    each $\xi(k,\theta,\beta_{k,\theta},s_{5})$ is an AH negative Einstein metric if $\beta_{k,\theta}=0$, an AC non-Einstein Ricci soliton if $\beta_{k,\theta}>0$.
Theorem 1.3.
For each $k\geq 3$ and $\theta\in(0,\pi)$, there exists an $\alpha_{k,\theta}\geq 0$ such that $\{\xi(k,\theta,s_{4},0)\mid s_{4}\geq\alpha_{k,\theta}\}$ is a continuous $1$-parameter family of complete steady Ricci solitons. In particular,
  1. (1)
    each $\xi(k,\theta,\alpha_{k,\theta},0)$ is an AC or an ALC Ricci-flat metric if $\alpha_{k,\theta}=0$;
  2. (2)
    each $\xi(k,\theta,\alpha_{k,\theta},0)$ is an AP non-Einstein steady soliton if $\alpha_{k,\theta}>0$. The base of the asymptotic paraboloid is the the Jensen $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$.
For comparison, we list previous examples and new Ricci solitons in Theorem 1.1-1.3 in the following table.
Metric Asymptotic Base Source
$\xi(k,0,0,0)$ almost-Hermitian Ricci-flat ALC Fubini–Study $\mathbb{CP}^{2m+1}/\mathbb{Z}_{k}$ [BB82]
$1\leq k\leq 2m+1$
$\xi(2m+2,0,0,0)$ almost-Kähler Ricci-flat AC standard $\mathbb{S}^{4m+3}/\mathbb{Z}_{2m+2}$
$\xi(k,0,0,s_{5})$ negative Einstein AH
$k\geq 1$, $s_{5}>0$
$\xi(2m+2,0,s_{4},0)$ steady Kähler–Ricci soliton ACP Fubini–Study $\mathbb{CP}^{2m+1}/\mathbb{Z}_{2m+2}$ [Cao96]
$s_{4}>0$
$\xi(k,0,s_{4},k-(2m+2))$ expanding Kähler–Ricci soliton AC [FIK03]
$k\geq 2m+3$, $s_{4}>0$
$\xi(k,0,s_{4},0)$ steady Ricci soliton ACP/AP Fubini–Study $\mathbb{CP}^{2m+1}/\mathbb{Z}_{k}$ [Sto24]
$k\geq 1$, $s_{4}>0$
$\xi(k,0,s_{4},0)$ steady Ricci soliton ACP Fubini–Study $\mathbb{CP}^{2m+1}/\mathbb{Z}_{k}$ [Win23]
$1\leq k\leq\sqrt{2m+2}$, $s_{4}>0$
$\xi(k,0,s_{4},s_{5})$ expanding Ricci soliton AC
$1\leq k\leq\sqrt{2m+2}$, $s_{4}>0$, $s_{5}>0$
$\xi(k,0,s_{4},0)$ steady Ricci soliton ACP Fubini–Study $\mathbb{CP}^{2m+1}/\mathbb{Z}_{k}$ [App17]
$1\leq k\leq 2m+2$, $s_{4}>0$
$\xi(k,0,s_{4},0)$ steady Ricci soliton ACP Fubini–Study $\mathbb{CP}^{2m+1}/\mathbb{Z}_{k}$
$k\geq 2m+3$, $s_{4}>\alpha_{k,0}$
$\xi(k,0,\alpha_{k,0},0)$ steady Ricci soliton AP standard $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$
$k\geq 2m+3$
$\xi(k,\theta,0,0)$ Ricci-flat ALC non-Kähler $\mathbb{CP}^{2m+1}/\mathbb{Z}_{k}$ Theorem 1.1
$\xi(k,\theta,0,s_{5})$ negative Einstein AH $k\in\{1,2\}$,
$\xi(k,\theta,s_{4},0)$ steady Ricci soliton ACP non-Kähler $\mathbb{CP}^{2m+1}/\mathbb{Z}_{k}$ $\theta\in(0,\pi)$,
$\xi(k,\theta,s_{4},s_{5})$ expanding Ricci soliton AC $s_{4}>0$, $s_{5}>0$
$\xi(k,\theta,s_{4},s_{5})$ expanding Ricci soliton AC Theorem 1.2
$\xi(k,\theta,\beta_{k,\theta},s_{5})$ if $\beta_{k,\theta}>0$: expanding Ricci soliton AC $k\geq 3$, $\theta\in(0,\pi)$,
if $\beta_{k,\theta}=0$: negative Einstein AH $s_{4}>\beta_{k,\theta}$, $s_{5}>0$
$\xi(k,\theta,\alpha_{k,\theta},0)$ if $\alpha_{k,\theta}=0$: Ricci-flat AC Jensen $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$ Theorem 1.3
ALC non-Kähler $\mathbb{CP}^{2m+1}$ $k\geq 3$,
$\xi(k,\theta,\alpha_{k,\theta},0)$ if $\alpha_{k,\theta}>0$: steady Ricci soliton AP Jensen $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$ $\theta\in(0,\pi)$,
$\xi(k,\theta,s_{4},0)$ steady Ricci soliton ACP non-Kähler $\mathbb{CP}^{2m+1}/\mathbb{Z}_{k}$ $s_{4}>\alpha_{k,\theta}$
Table 1.
Compared to Theorem 1.1, the asymptotics for steady solitons in Theorem 1.3 are less settled, as they depend on the value of $\alpha_{k,\theta}$. The boundary cases where $\theta\in\{0,\pi\}$, however, are better understood in the previous examples. Specifically, the metrics $\xi(2m+2,0,0,0)$ and $\xi(2,\pi,0,0)$ are AC Ricci-flat metrics, with the asymptotic cones’ bases being the standard orbifolds $\mathbb{S}^{4m+3}/\mathbb{Z}_{2m+2}$ and $\mathbb{S}^{3}/\mathbb{Z}_{2}$, respectively; see [BB82]. The existence of $\alpha_{k,0}$ and $\alpha_{k,\pi}$ are established in [App17, Theorem 5.1]. In particular, we have $\alpha_{k,0}>0$ if $k\geq 2m+3$, and $\alpha_{k,\pi}>0$ if $k\geq 3$. Thanks to these results, we can further obtain the following existence theorem for some non-collapsed steady Ricci solitons.
Theorem 1.4.
On each $\mathcal{O}(k)$ with $3\leq k\leq 2m+1$, there exists at least one AC Ricci-flat metric, whose base of the cone is the Jensen $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$. On each $\mathcal{O}(k)$ with $k\geq 3$, there exist infinitely many AP steady Ricci solitons whose base of the paraboloid is the Jensen $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$.
This article is structured as the following. We begin by presenting the cohomogeneity one Ricci soliton equation and introducing a new coordinate transformation, differing from that in [Chi26], to establish local existence. For global analysis, we first construct an invariant compact set to prove Theorem 1.1. Next, we generalize [App17, Theorem 5.1] and prove Theorem 1.2-1.3. Finally, we investigate the asymptotics for these Ricci solitons and prove Theorem 1.4.
2. Cohomogeneity one equation
The geometric setup is similar to the one in [Chi26]. Specifically, let $Q$ be the inner product on $\mathfrak{g}$ that generates the round metric on $\mathbb{S}^{4m+3}$. We have a $Q$-orthogonal decomposition of the isotropy representation $\mathfrak{g}/\mathfrak{k}$ into three irreducible summands. Consider the ansatz
(2.1) $g=dt^{2}+a^{2}(t)\left.Q\right|_{\mathbf{1}}+b^{2}(t)\left.Q\right|_{\mathbf{2}}+c^{2}(t)\left.Q\right|_{\mathbf{4m}},$
where $b$ and $c$ are metric components for the base space $\mathbb{CP}^{2m+1}$ and $a$ is for the $\mathbb{S}^{1}$-fiber. Since $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$ is locally isometric to $\mathbb{S}^{4m+3}$, the Ricci tensor of $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$ shares the same formula as the one of $\mathbb{S}^{4m+3}$. The equation for non-shrinking cohomogeneity one Ricci solitons is
(2.2) $\begin{split}\frac{\ddot{a}}{a}-\left(\frac{\dot{a}}{a}\right)^{2}&=-\left(\frac{\dot{a}}{a}+2\frac{\dot{b}}{b}+4m\frac{\dot{c}}{c}\right)\frac{\dot{a}}{a}+2\frac{a^{2}}{b^{4}}+4m\frac{a^{2}}{c^{4}}+\frac{\dot{a}}{a}\dot{f}+\frac{\epsilon}{2},\\ \frac{\ddot{b}}{b}-\left(\frac{\dot{b}}{b}\right)^{2}&=-\left(\frac{\dot{a}}{a}+2\frac{\dot{b}}{b}+4m\frac{\dot{c}}{c}\right)\frac{\dot{b}}{b}+\frac{4}{b^{2}}-2\frac{a^{2}}{b^{4}}+4m\frac{b^{2}}{c^{4}}+\frac{\dot{b}}{b}\dot{f}+\frac{\epsilon}{2},\\ \frac{\ddot{c}}{c}-\left(\frac{\dot{c}}{c}\right)^{2}&=-\left(\frac{\dot{a}}{a}+2\frac{\dot{b}}{b}+4m\frac{\dot{c}}{c}\right)\frac{\dot{c}}{c}+\frac{4m+8}{c^{2}}-2\frac{a^{2}}{c^{4}}-4\frac{b^{2}}{c^{4}}+\frac{\dot{c}}{c}\dot{f}+\frac{\epsilon}{2},\\ \ddot{f}+\frac{\epsilon}{2}&=\frac{\ddot{a}}{a}+2\frac{\ddot{b}}{b}+4m\frac{\ddot{c}}{c}.\end{split}$
By [Ham93, Theorem 20.1], the equality
(2.3) $\ddot{f}+(\mathrm{tr}(L)-\dot{f})\dot{f}-\epsilon f=C$
holds for some constant $C$. Canceling all second order terms using (2.2) and (2.3), we obtain the following equation:
(2.4) $\begin{split}&\mathrm{tr}(L^{2})+r_{s}+\frac{n-1}{2}\epsilon-(\dot{f}-\mathrm{tr}(L))^{2}=C+\epsilon f.\end{split}$
By [Che09, Corollary 2.3], it can be deduced that $C+\epsilon f<0$ for a non-Einstein Ricci soliton.
We set $f(0)=0$ without loss of generality. By [Buz11], the potential function $f$ is an even function in $t$ near 0. As the $\mathbb{S}^{1}$-fiber collapses at the singular orbit, we have $\mathrm{tr}(L)=\frac{1}{t}+O(t)$ as $t\to 0$. Therefore, we have $2\ddot{f}(0)=C$ by(2.3). As the radius for the $\mathbb{S}^{1}$-fiber of $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$ is $\frac{1}{k}$, we require $\dot{a}(0)=k$ for smooth extension and we consider the initial condition
(2.5) $\lim_{t\to 0}(a,b,c,f,\dot{a},\dot{b},\dot{c})=\left(0,b_{0},c_{0},0,k,0,0\right),\quad b_{0},c_{0}>0.$
With each fixed $k$, parameters $b_{0}$, $c_{0}$ and $C$ are free. The equation (2.2) is invariant under homothety change if $\epsilon=0$. Therefore, the above initial condition has two free parameters for the steady case. Note that the singular orbit is not round if $b_{0}\neq c_{0}$.
Apply coordinate change $d\eta=(\mathrm{tr}(L)-\dot{f})dt$. Define
(2.6) $\begin{split}&X_{1}=\frac{\frac{\dot{a}}{a}}{\mathrm{tr}(L)-\dot{f}},\quad X_{2}=\frac{\frac{\dot{b}}{b}}{\mathrm{tr}(L)-\dot{f}},\quad X_{3}=\frac{\frac{\dot{c}}{c}}{\mathrm{tr}(L)-\dot{f}},\\ &Z_{1}=\frac{a^{2}}{b^{2}},\quad Z_{2}=\frac{\frac{1}{b^{2}}}{(\mathrm{tr}(L)-\dot{f})^{2}},\quad Z_{3}=\frac{\frac{b^{2}}{c^{4}}}{(\mathrm{tr}(L)-\dot{f})^{2}},\quad Z_{4}=\frac{\frac{1}{c^{2}}}{(\mathrm{tr}(L)-\dot{f})^{2}},\quad W=\frac{1}{(\mathrm{tr}(L)-\dot{f})^{2}}.\end{split}$
Define functions on $\eta$
(2.7) $\begin{split}&R_{1}=2Z_{1}Z_{2}+4mZ_{1}Z_{3},\\ &R_{2}=4Z_{2}-2Z_{1}Z_{2}+4mZ_{3},\\ &R_{3}=(4m+8)Z_{4}-2Z_{1}Z_{3}-4Z_{3},\\ &R_{s}=R_{1}+2R_{2}+4mR_{3}=8Z_{2}+4m(4m+8)Z_{4}-8mZ_{3}-2Z_{1}Z_{2}-4mZ_{1}Z_{3},\\ &H=X_{1}+2X_{2}+4mX_{3},\quad G=X_{1}^{2}+2X_{2}^{2}+4mX_{3}^{2}.\end{split}$
Let denote the derivative with respect to $\eta$. The soliton equation (2.2) becomes a polynomial system
(2.8) $\begin{bmatrix}X_{1}\\ X_{2}\\ X_{3}\\ Z_{1}\\ Z_{2}\\ Z_{3}\\ Z_{4}\\ W\end{bmatrix}^{\prime}=V(X_{1},X_{2},X_{3},Y_{1},Y_{2},Y_{3},W):=\begin{bmatrix}V_{1}\\ V_{2}\\ V_{3}\\ V_{4}\\ V_{5}\\ V_{6}\\ V_{7}\\ V_{8}\end{bmatrix}=\begin{bmatrix}X_{1}(G-\frac{\epsilon}{2}W-1)+R_{1}+\frac{\epsilon}{2}W\\ X_{2}(G-\frac{\epsilon}{2}W-1)+R_{2}+\frac{\epsilon}{2}W\\ X_{3}(G-\frac{\epsilon}{2}W-1)+R_{3}+\frac{\epsilon}{2}W\\ 2Z_{1}(X_{1}-X_{2})\\ 2Z_{2}(G-\frac{\epsilon}{2}W-X_{2})\\ 2Z_{3}(G-\frac{\epsilon}{2}W+X_{2}-2X_{3})\\ 2Z_{4}(G-\frac{\epsilon}{2}W-X_{3})\\ 2W(G-\frac{\epsilon}{2}W)\end{bmatrix}$
Define the polynomial $Q:=G+R_{s}+(n-1)\frac{\epsilon}{2}W-1$. The conservation law (2.4) becomes
(2.9) $\begin{split}\frac{C+\epsilon f}{(\mathrm{tr}(L)-\dot{f})^{2}}=Q.\end{split}$
As discussed above, the quantity $C+\epsilon f$ (and hence $Q$) remains negative for a non-Einstein Ricci soliton. We focus on the following $7$-dimensional algebraic set.
$\mathcal{RS}:=\{Q\leq 0\}\cap\{H\leq 1\}\cap\{W\geq 0\}\cap\{Z_{1},Z_{2},Z_{3},Z_{4}\geq 0\}\cap\{Z_{4}^{2}=Z_{2}Z_{3}\}.$
Since
[…]
it is clear that $\mathcal{RS}$ is flow-invariant, a slight generalization to [DW09, Proposition 3.1]. For the expanding case, we fix the homothety by setting $\epsilon=1$. For $\epsilon=0$, the homothety change is transformed to translation on the variable $\eta$ in the new coordinate. Therefore, different trajectories to system (2.8) in $\mathcal{RS}$ represent different homothety families of Ricci solitons.
In [Chi26], we transform (2.2) using the $XYW$-coordinates, where
(2.11) $Y_{1}=\sqrt{Z_{1}},\quad Y_{2}=\sqrt{Z_{2}},\quad Y_{3}=\sqrt{Z_{3}}.$
Many computations in [Chi26] carry over via (2.11). In particular, the original coordinates are recovered with the same formula. Local solutions that satisfy the initial condition (2.5) is transformed integral curves that emanate $p_{0}:=(1,0,0,0,0,0,0,0,0)$. The positiviteness and finiteness of $b_{0}$ and $c_{0}$ requires the limit
(2.12) $\lim\limits_{\eta\to-\infty}\sqrt{\frac{Z_{3}}{Z_{2}}}>0$
to exist. If $\epsilon=1$, the zeroth order initial condition $\lim\limits_{t\to 0}(b,c)=(b_{0},c_{0})$ becomes
(2.13) $\lim\limits_{\eta\to-\infty}\frac{W}{Z_{2}}=b_{0}^{2},\quad\lim\limits_{\eta\to-\infty}\frac{W}{Z_{3}}=\frac{c_{0}^{2}}{b_{0}},\quad\lim\limits_{\eta\to-\infty}\frac{W}{Z_{4}}=c_{0}^{2}.$
The first order initial condition $\lim\limits_{t\to 0}(\dot{a},\dot{b},\dot{c})=(k,0,0)$ becomes
(2.14) $\lim\limits_{\eta\to-\infty}\sqrt{\frac{Z_{1}}{Z_{2}}}=k,\quad\lim\limits_{\eta\to-\infty}\frac{X_{2}}{\sqrt{Z_{2}}}=0,\quad\lim\limits_{\eta\to-\infty}\frac{X_{3}}{\sqrt{Z_{4}}}=0.$
The problem of finding complete Ricci solitons are transformed to finding integral curves emanate $p_{0}$ that are defined on $\mathbb{R}$; see [Chi26, Proposition 2.2].
We end this section by listing some important invariant subsets of lower dimensions below.
  1. (1)
    $\mathcal{RS}_{\text{Einstein}}:=\mathcal{RS}\cap\{Q=0\}\cap\{H=1\}$
    As discussed above, equalities $Q=0$ and $H=1$ mean that $f$ is identically zero and $C=0$. The subsystem restricted on $\mathcal{E}$ is hence the cohomogeneity one non-positive Einstein equation.
  2. (2)
    $\mathcal{RS}_{\text{steady}}:=\mathcal{RS}\cap\{W=0\}$
    The variable $W$ is decoupled if $\epsilon=0$. Hence a solution to the subsystem restricted on $\mathcal{RS}_{\text{steady}}$ represents a steady Ricci soliton even though the function $\frac{1}{(\mathrm{tr}(L)-\dot{f})^{2}}$ does not identically vanish.
  3. (3)
    $\mathcal{RF}:=\mathcal{RS}_{\text{steady}}\cap\mathcal{RS}_{\text{Einstein}}$
    The subsystem restricted on $\mathcal{RF}$ is the cohomogeneity one Ricci-flat equation.
  4. (4)
    $\mathcal{RS}_{\text{FS}}:=\mathcal{RS}\cap\{Z_{2}-Z_{3}=0\}\cap\{X_{3}-X_{3}=0\}$
    Trajectories on this set represent metrics whose singular orbit is the Fubini–Study $\mathbb{CP}^{2m+1}$. The Kähler class of $\mathbb{CP}^{2m+1}$ is $k$ multiple of indivisible integral cohomology class in $H^{2}(\mathbb{CP}^{2m+1},\mathbb{Z})$. For integral curves that are not in $\mathcal{RS}_{\text{FS}}$, the base space is not Kähler.
  5. (5)
    $\mathcal{RS}_{\text{KE}}:=\mathcal{RS}_{\text{FS}}\cap\{X_{2}^{2}=Z_{1}Z_{2}\}\cap\{(4m+4)Z_{2}+\frac{\epsilon}{2}W-X_{2}(1+X_{1})=0\}$
    Trajectories on this set represent Kähler–Ricci solitons; see [DW11]. Note that the base space must be the Fubini–Study $\mathbb{CP}^{2m+1}$.
  6. (6)
    $\mathcal{RS}_{\text{round}}:=\mathcal{RS}\cap\{Z_{1}=1\}\cap\{X_{1}-X_{2}=0\}$
    Trajectories on this set represent metrics with a principal orbit as a round $\mathbb{S}^{3}$-bundle over $\mathbb{HP}^{m}$.
  7. (7)
    $\mathcal{RS}_{m=0}:=\mathcal{RS}_{\text{steady}}\cap\{X_{3},Z_{3}=0\}$.
    Although integral curves in $\mathcal{RS}_{m=0}$ are geometrically degenerate when viewed as $(4m+4)$-dimensional Ricci solitons, they can be interpreted as four-dimensional steady Ricci solitons on complex line bundles over $\mathbb{CP}^{1}$. Indeed, setting $m=\epsilon=0$ in (2.2) and ignoring the function $c$ yields precisely the system for four-dimensional steady Ricci solitons as in [App17, (2.1)–(2.3)]. Applying the coordinate transformation (2.6), this system is equivalent to the restriction of (2.8) to $\mathcal{RS}_{\text{steady}}\cap\mathcal{RS}_{\text{FS}}$ with $m=0$, which in turn coincides with the dynamical system on $\mathcal{RS}_{m=0}$. It is important to note that, unlike the four-dimensional expanders studied in [Wan25, Don25], this $m=0$ subsystem produces only steady solitons.
3. Local Analysis
The linearization at $p_{0}$ is
(3.1) $\begin{bmatrix}2&0&0&0&0&0&0&0\\ 0&0&0&0&4&4m&0&\frac{\epsilon}{2}\\ 0&0&0&0&0&-4&4m+8&\frac{\epsilon}{2}\\ 0&0&0&2&0&0&0&0\\ 0&0&0&0&2&0&0&0\\ 0&0&0&0&0&2&0&0\\ 0&0&0&0&0&0&2&0\\ 0&0&0&0&0&0&0&2\\ \end{bmatrix}$.
The linearization has two zero eigenvalues. The other eigenvalue of $2$ has multiplicity of $6$. Unstable eigenvectors are:
$w_{1}=\begin{bmatrix}-(4m+2)(2m+2)\\ 2m+2\\ 2m+2\\ 0\\ 1\\ 1\\ 1\\ 0\end{bmatrix},w_{2}=\begin{bmatrix}-4\\ 2\\ 0\\ 0\\ 1\\ 0\\ 0\\ 0\end{bmatrix},w_{3}=\begin{bmatrix}-4(m+1)^{2}\sqrt{2}\\ 2\sqrt{2}\\ (m+2)\sqrt{2}\\ 0\\ \sqrt{2}\\ 0\\ \frac{\sqrt{2}}{2}\\ 0\end{bmatrix},w_{4}=\begin{bmatrix}-1\\ 0\\ 0\\ 0\\ 0\\ 0\\ 0\\ 0\end{bmatrix},w_{5}=\begin{bmatrix}-(4m+2)\frac{\epsilon}{2}\\ \frac{\epsilon}{2}\\ \frac{\epsilon}{2}\\ 0\\ 0\\ 0\\ 0\\ 2\end{bmatrix},w_{6}=\begin{bmatrix}0\\ 0\\ 0\\ 1\\ 0\\ 0\\ 0\\ 0\end{bmatrix}$.
Remark 3.1.
We apply the XZWXZW-coordinates instead of the XYWXYW-coordinates in [Chi26] because the linearization at $p_{0}$ has equal positive eigenvalues. This property allows us to apply [CL55, Chapter 13, Theorem 4.5] to analyze functions along integral curves near $p_{0}$.
Since $p_{0}$ is not hyperbolic, the Hartman–Grobman Theorem does not directly apply. We apply center manifold theory for the local analysis near $p_{0}$; see [Car12] and [Per13]. Consider the linear transformation
(3.2) $\tilde{X}_{1}=X_{1}-1,\quad B:=X_{2}-2Z_{2}-2mZ_{3}-\frac{\epsilon}{4}W,\quad C:=X_{3}+2Z_{3}-(2m+4)Z_{4}-\frac{\epsilon}{4}W.$
The dynamical system (2.8) is topologically conjugate to
(3.3) $\begin{bmatrix}\tilde{X}_{1}\\ B\\ C\\ Z_{1}\\ Z_{2}\\ Z_{3}\\ Z_{4}\\ W\end{bmatrix}^{\prime}=F(\tilde{X}_{1},B,C,Z_{1},Z_{2},Z_{3},Z_{4},W)=\begin{bmatrix}V_{1}\\ F_{B}\\ F_{C}\\ V_{4}\\ V_{5}\\ V_{6}\\ V_{7}\\ V_{8}\end{bmatrix}=\begin{bmatrix}V_{1}\\ V_{2}-2V_{5}-2mV_{6}-\frac{\epsilon}{4}V_{8}\\ V_{3}+2V_{6}-(2m+4)V_{7}-\frac{\epsilon}{4}V_{8}\\ V_{4}\\ V_{5}\\ V_{6}\\ V_{7}\\ V_{8}\end{bmatrix}$
where $V_{i}$ are functions in (2.8) with $X_{i}$ being functions defined by (3.2). The linearization of (3.3) at $p_{0}$ is
(3.4) $\begin{bmatrix}2&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&2&0&0&0&0\\ 0&0&0&0&2&0&0&0\\ 0&0&0&0&0&2&0&0\\ 0&0&0&0&0&0&2&0\\ 0&0&0&0&0&0&0&2\\ \end{bmatrix}$.
By [Per13, Theorem 2, Page 161], the dynamical system (3.3) is topologically conjugate to
(3.5) $\begin{bmatrix}\tilde{X}_{1}\\ B\\ C\\ Z_{1}\\ Z_{2}\\ Z_{3}\\ Z_{4}\\ W\end{bmatrix}^{\prime}=\begin{bmatrix}2\tilde{X}_{1}\\ F_{B}(\tilde{x}_{1},B,C,z_{1},z_{2},z_{3},z_{4},w)\\ F_{C}(\tilde{x}_{1},B,C,z_{1},z_{2},z_{3},z_{4},w)\\ 2Z_{1}\\ 2Z_{2}\\ 2Z_{3}\\ 2Z_{4}\\ 2W\end{bmatrix}$,
where lower case letters are functions of $(B,C)$ that satisfy
(3.6) $\begin{bmatrix}\partial_{B}\tilde{x}_{1}&\partial_{C}\tilde{x}_{1}\\ \partial_{B}z_{1}&\partial_{C}z_{1}\\ \partial_{B}z_{2}&\partial_{C}z_{2}\\ \partial_{B}z_{3}&\partial_{C}z_{3}\\ \partial_{B}z_{4}&\partial_{C}z_{4}\\ \partial_{B}w&\partial_{C}w\end{bmatrix}\begin{bmatrix}F_{B}(\tilde{x}_{1},B,C,z_{1},z_{2},z_{3},z_{4},w)\\ F_{C}(\tilde{x}_{1},B,C,z_{1},z_{2},z_{3},z_{4},w)\end{bmatrix}=\begin{bmatrix}V_{1}(\tilde{x}_{1},B,C,z_{1},z_{2},z_{3},z_{4},w)\\ V_{4}(\tilde{x}_{1},B,C,z_{1},z_{2},z_{3},z_{4},w)\\ V_{5}(\tilde{x}_{1},B,C,z_{1},z_{2},z_{3},z_{4},w)\\ V_{6}(\tilde{x}_{1},B,C,z_{1},z_{2},z_{3},z_{4},w)\\ V_{7}(\tilde{x}_{1},B,C,z_{1},z_{2},z_{3},z_{4},w)\\ V_{8}(\tilde{x}_{1},B,C,z_{1},z_{2},z_{3},z_{4},w)\end{bmatrix}$
(3.7) $\begin{bmatrix}\tilde{x}_{1}\\ z_{1}\\ z_{2}\\ z_{3}\\ z_{4}\\ w\end{bmatrix}(0,0)=\begin{bmatrix}0\\ 0\\ 0\\ 0\\ 0\\ 0\\ \end{bmatrix},\quad\begin{bmatrix}\partial_{B}\tilde{x}_{1}&\partial_{C}\tilde{x}_{1}\\ \partial_{B}z_{1}&\partial_{C}z_{1}\\ \partial_{B}z_{2}&\partial_{C}z_{2}\\ \partial_{B}z_{3}&\partial_{C}z_{3}\\ \partial_{B}z_{4}&\partial_{C}z_{4}\\ \partial_{B}w&\partial_{C}w\end{bmatrix}(0,0)=\begin{bmatrix}0&0\\ 0&0\\ 0&0\\ 0&0\\ 0&0\\ 0&0\\ \end{bmatrix}$
and can be approximated by power series in $(B,C)$.
With functions $F_{B}$ and $F_{C}$ explicitly known in (3.3), we claim that
(3.8) $z_{2}=0,z_{3}=0,z_{4}=0,w=0$
as power series of $(B,C)$. Suppose otherwise, we have
$\begin{bmatrix}F_{B}\\ F_{C}\end{bmatrix}\sim\begin{bmatrix}4z_{2}+4mz_{3}+\frac{\epsilon}{2}w\\ (4m+8)z_{4}-4z_{3}+\frac{\epsilon}{2}w\end{bmatrix},\quad\begin{bmatrix}V_{5}\\ V_{6}\\ V_{7}\\ V_{8}\end{bmatrix}\sim\begin{bmatrix}2z_{2}\\ 2z_{3}\\ 2z_{4}\\ 2w\end{bmatrix}$
By the last four lines in (3.6), the lowest power on the left hand side is strictly larger than the lowest power on the right hand side. Therefore, we must have power series (3.8). Furthermore, we have
$\begin{bmatrix}F_{B}\\ F_{C}\end{bmatrix}\sim\begin{bmatrix}2\tilde{x}_{1}B\\ 2\tilde{x}_{1}C\end{bmatrix},\quad\begin{bmatrix}V_{1}\\ V_{4}\end{bmatrix}\sim\begin{bmatrix}2\tilde{x}_{1}\\ 2z_{1}\end{bmatrix}$
Note that the lowest powers of $\tilde{x}_{1}B$ and $\tilde{x}_{1}C$ are at least $3$. From the first two lines in (3.6), we further conclude that $\tilde{x}_{1}=0$ and $z_{1}=0$ as power series of $(B,C)$. Therefore, the dynamical system (2.8) is topologically conjugate to
(3.9) $\begin{bmatrix}\tilde{X}_{1}\\ B\\ C\\ Z_{1}\\ Z_{2}\\ Z_{3}\\ Z_{4}\\ W\end{bmatrix}^{\prime}=\begin{bmatrix}2\tilde{X}_{1}\\ B(2B^{2}+4mC^{2})\\ C(2B^{2}+4mC^{2})\\ 2Z_{1}\\ 2Z_{2}\\ 2Z_{3}\\ 2Z_{4}\\ 2W\end{bmatrix}$,
whose general solutions are
(3.10) $\begin{bmatrix}\tilde{X}_{1}\\ B\\ C\\ Z_{1}\\ Z_{2}\\ Z_{3}\\ Z_{4}\\ W\end{bmatrix}=\begin{bmatrix}u_{1}e^{2\eta}\\ B_{0}\frac{1}{\sqrt{a_{1}-\eta}}\\ C_{0}\frac{1}{\sqrt{a_{2}-\eta}}\\ u_{4}e^{2\eta}\\ u_{5}e^{2\eta}\\ u_{6}e^{2\eta}\\ u_{7}e^{2\eta}\\ u_{8}e^{2\eta}\\ \end{bmatrix}$.
In particular, integral curves with non-zero $B_{0}$ or $C_{0}$ are center manifolds of $p_{0}$. These integral curves are geometrically degenerate. Specifically, by (2.14) and the existence of (2.13), we have
(3.11) $\begin{split}\lim\limits_{\eta\to-\infty}\frac{B}{\sqrt{Z_{2}}}&=\lim\limits_{\eta\to-\infty}\frac{X_{2}-2Z_{2}-2mZ_{3}-\frac{\epsilon}{4}W}{\sqrt{Z}_{2}}=0,\\ \lim\limits_{\eta\to-\infty}\frac{C}{\sqrt{Z_{4}}}&=\lim\limits_{\eta\to-\infty}\frac{X_{3}+2Z_{3}-(2m+4)Z_{4}-\frac{\epsilon}{4}W}{\sqrt{Z_{4}}}=0.\end{split}$
By (3.10), the above limits hold only if $B_{0}=C_{0}=0$. Therefore, we only consider integral curves in the unstable manifold of $p_{0}$. By [CL55, Chapter 13, Theorem 4.5] and the analysis above, there is no ambiguity to let $\xi(s_{1},s_{2},s_{3},s_{4},s_{5},s_{6})$ denote the integral curve of (2.8), where
(3.12) $\begin{split}&\xi(s_{1},s_{2},s_{3},s_{4},s_{5},s_{6})\\ &=p_{0}+s_{1}e^{2\eta}w_{1}+s_{2}e^{2\eta}w_{2}+s_{3}e^{2\eta}w_{3}+s_{4}e^{2\eta}w_{4}+s_{5}e^{2\eta}w_{5}+s_{6}e^{2\eta}w_{6}+O(e^{(2+\delta)\eta}).\end{split}$
The first three parameters control how the principal orbit is squashed. We fix $s_{1}^{2}+s_{2}^{2}+s_{3}^{2}=1$ to mod out the scalar multiplication of the vector $\sum_{i=1}^{6}s_{i}w_{i}$. As $p_{0}$ is at the tip of the constraint cone $\{Z_{4}^{2}=Z_{2}Z_{3}\}$, the normalized initial velocity must be tangent to the cone. Since
(3.13) $Z_{2}\sim\left(s_{1}+s_{2}+\sqrt{2}s_{3}\right)e^{2\eta},\quad Z_{3}\sim s_{1}e^{2\eta},\quad Z_{4}\sim\left(s_{1}+\frac{\sqrt{2}}{2}s_{3}\right)e^{2\eta}$
near $p_{0}$, together with the constraint $Z_{2}Z_{3}=Z_{4}^{2}$, we have
(3.14) $s_{1}+s_{2}+\sqrt{2}s_{3}\geq 0,\quad s_{1}\geq 0,\quad s_{1}+\frac{\sqrt{2}}{2}s_{3}\geq 0,\quad s_{3}^{2}=2s_{1}s_{2}.$
We further impose the condition that $s_{2}+2s_{3}\geq 0$ so that $Z_{2}-Z_{3}\geq 0$ initially. Combining all the above conditions, the parameters $(s_{1},s_{2},s_{3})$ are chosen from the circular arc
(3.15) $\begin{split}&\left\{(s_{1},s_{2},s_{3})\mid s_{1}^{2}+s_{2}^{2}+s_{3}^{2}=1,\quad s_{3}^{2}=2s_{1}s_{2},\quad s_{1},s_{2},s_{3}\geq 0\right\}\\ &=\left\{(s_{1},s_{2},s_{3})\mid(s_{1}-s_{2})^{2}+2s_{3}^{2}=1,\quad s_{1}+s_{2}=1,\quad s_{3}\geq 0\right\}.\end{split}$
On the other hand, the smoothness condition requires $\lim\limits_{t\to 0}\dot{a}=k$. By (2.14), we must set $s_{6}=k^{2}(s_{1}+s_{2}+\sqrt{2}s_{3}).$ Parametrize (3.15) by
(3.16) $s_{1}=\frac{1+\cos(\theta)}{2},\quad s_{2}=\frac{1-\cos(\theta)}{2},\quad s_{3}=\frac{1}{\sqrt{2}}\sin(\theta),\quad\theta\in[0,\pi].$
We write
$w(\theta,k)=\frac{1+\cos(\theta)}{2}w_{1}+\frac{1-\cos(\theta)}{2}w_{2}+\frac{1}{\sqrt{2}}\sin(\theta)w_{3}+k^{2}\left(1+\sin(\theta)\right)w_{6}.$
All unstable eigenvectors are perpendicular to $(\nabla W)(p_{0})$ except $w_{4}$. Steady Ricci solitons are represented by integral curves with $s_{5}=0$. Expanding Ricci solitons are represented by integral curves with $s_{5}>0$. A straightforward computation shows that
$(\nabla Q)(p_{0})=\begin{bmatrix}2\\ 0\\ 0\\ 0\\ 8\\ -8m\\ 4m(4m+8)\\ (n-1)\frac{\epsilon}{2}\end{bmatrix}$.
Note that $w_{4}$ is the only unstable eigenvectors that is not perpendicular to $(\nabla Q)(P_{\mathbb{CP}^{2m+1}})$ or $(\nabla H)(P_{\mathbb{CP}^{2m+1}})$. By (3.19), we have
(3.17) $Q\sim-2s_{4}e^{2\eta},\quad 1-H\sim s_{4}e^{2\eta}.$
Therefore,
(3.18) $\frac{\ddot{f}}{(\mathrm{tr}(L)-\dot{f})^{2}}=Q-H+1\sim-s_{4}e^{2\eta}$
as $\eta\to$-$\infty$. The parameter $s_{4}$ controls the initial condition $\ddot{f}(0)$ and we set $s_{4}\geq 0$. If integral curves with $s_{4}$ represent cohomogeneity one Einstein metrics.
In summary, we consider the following $4$-parameter family of integral curve such that
(3.19) $\begin{split}\xi(k,\theta,s_{4},s_{5})=p_{0}+e^{2\eta}w(\theta,k)+s_{4}e^{2\eta}w_{4}+s_{5}e^{2\eta}w_{5}+O(e^{(2+\delta)\eta}),\quad\theta\in[0,\pi],k\in\mathbb{Z}^{+},s_{4},s_{5}\geq 0.\end{split}$
For $\theta\in[0,\pi)$, these integral curves are identified as local solutions on a tubular neighborhood around $\mathbb{CP}^{2m+1}$. For $\theta=\pi$, the integral curves are in $\mathcal{RS}_{m=0}$, and they represent local solutions near $\mathbb{CP}^{1}$. The extra parameter $k$ creates the illusion of an additional degree of freedom to extend the local solution smoothly to the base space. However, for each fixed $\mathcal{O}(k)$, the smoothness condition only allows varying three parameters from $(\theta,s_{4},s_{5})$, as predicted in [EW00].
Remark 3.2.
Unlike the 1-parameter family of Kähler–Ricci solitons on $\mathbb{C}^{2m+2}$ studied in [Cao97], the Ricci solitons represented by (3.19) have mixed signs in their sectional curvatures. Specifically, consider the normal-tangential sectional curvatures $K_{i}:=X_{i}-X_{i}^{2}-R_{i}-\frac{\epsilon}{2}W$. We observe that $K_{1}>0$, while $K_{2},K_{3}<0$ initially along each $\xi(k,\theta,s_{4},s_{5})$. The non-negativity of sectional curvatures significantly constrains the topology of steady gradient Ricci solitons; for a detailed discussion, see the recent developments in [DH24].
4. Global analysis
In this section, we first construct a compact invariant set $\mathcal{F}$ to prove the completeness of Ricci solitons in Theorem 1.1; see (4.5). Each $\xi(k,\theta,s_{4},s_{5})$ with $k\geq 3$ is not in $\mathcal{F}$ initially. Therefore, we generalize [App17, Theorem 5.1] to establish the existence of $\beta_{\theta,k}$ and $\alpha_{\theta,k}$ in Theorem 1.2-1.3. By the continuous dependence of parameters, we also establish the existence of complete non-collapsed steady Ricci solitons in Theorem 1.4.
Define the function
$F_{l}:=X_{2}-X_{1}+l\left(\sqrt{\frac{Z_{2}}{Z_{1}}}-\sqrt{Z_{1}Z_{2}}\right)$
for some constant $l$. We have
(4.1) $\begin{split}\langle\nabla F_{l},V\rangle&=F_{l}\left(G-\frac{\epsilon}{2}W-1+2l\sqrt{Z_{1}Z_{2}}\right)\\ &\quad+\left(l\sqrt{\frac{Z_{2}}{Z_{1}}}(1-X_{1})+(4-2l^{2})Z_{2}+4mZ_{3}\right)(1-Z_{1}).\end{split}$
On $\mathcal{RS}\cap\{F_{l}=0\}\cap\{Z_{2}-Z_{3}\geq 0\}$, we have
(4.2) $\begin{split}1&\geq G+R_{s}\\ &=X_{1}^{2}+2\left(X_{1}-l\left(\sqrt{\frac{Z_{2}}{Z_{1}}}-\sqrt{Z_{1}Z_{2}}\right)\right)^{2}+4mX_{3}^{2}\\ &\quad+8Z_{2}+4m(4m+8)Z_{4}-8mZ_{3}-2Z_{1}Z_{2}-4mZ_{1}Z_{3}\end{split}$
Proposition 4.1.
The inequality $\left.\langle\nabla F_{2},V\rangle\right|_{\{F_{2}=0\}}\geq 0$ holds on
$\mathcal{RS}\cap\{1-Z_{1}\geq 0\}\cap\{F_{2}=0\}\cap\{Z_{2}-Z_{3}\geq 0\}\cap\{X_{1}\geq 0\}$
Proof.
By inequalities $1-Z_{1}\geq 0$ and $Z_{2}-Z_{3}\geq 0$, the inequality (4.2) with $l=2$ becomes
(4.3) $\begin{split}1&\geq X_{1}^{2}+2\left(X_{1}-2\left(\sqrt{\frac{Z_{2}}{Z_{1}}}-\sqrt{Z_{1}Z_{2}}\right)\right)^{2}+8Z_{2}-2Z_{1}Z_{2}\\ &=2\left(X_{1}-2\sqrt{\frac{Z_{2}}{Z_{1}}}+\sqrt{Z_{1}Z_{2}}\right)^{2}+(X_{1}+2\sqrt{Z_{1}Z_{2}})^{2}\\ &\geq(X_{1}+2\sqrt{Z_{1}Z_{2}})^{2}.\end{split}$
Therefore, the inequality $1\geq X_{1}+2\sqrt{Z_{1}Z_{2}}$ holds on $\mathcal{RS}\cap\{1-Z_{1}\geq 0\}\cap\{F_{2}=0\}\cap\{Z_{2}-Z_{3}\geq 0\}\cap\{X_{1}\geq 0\}$. By (4.1), we have
(4.4) $\begin{split}\left.\langle\nabla F_{2},V\rangle\right|_{\{F_{2}=0\}}&=\left(2\sqrt{\frac{Z_{2}}{Z_{1}}}(1-X_{1})-4Z_{2}+4mZ_{3}\right)(1-Z_{1})\\ &\geq\left(1-X_{1}-2\sqrt{Z_{1}Z_{2}}\right)2\sqrt{\frac{Z_{2}}{Z_{1}}}(1-Z_{1})\\ &\geq 0.\end{split}$
The proof is complete. ∎
4.1. $k\leq 2$
We proceed to prove that each $\xi(\theta,k,s_{4},s_{5})$ with $k\in\{1,2\}$ is defined on $\mathbb{R}$. Define
(4.5) $\begin{split}\mathcal{F}&:=\mathcal{RS}\cap\{1-Z_{1}\geq 0\}\cap\left\{F_{2}\geq 0\right\}\\ &\quad\cap\{Z_{2}-Z_{3}\geq 0\}\cap\left\{2(\sqrt{Z_{2}}-\sqrt{Z_{3}})+X_{3}-X_{2}\geq 0\right\}\cap\{X_{1},X_{2},X_{3}\geq 0\}.\end{split}$
By [Chi26], we have the following lemma.
Lemma 4.2.
The set $\mathcal{F}$ is compact and invariant.
Proof.
With inequalities $1-Z_{1}\geq 0$ and $Z_{2}-Z_{3}\geq 0$, it is clear that $\mathcal{F}$ is compact. Furthermore, we have $R_{2},R_{3}\geq 0$. Therefore, the vector field $V$ restricted on faces $\mathcal{F}\cap\left\{X_{i}=0\right\}$ points inward.
For the face $\mathcal{F}\cap\{Z_{2}-Z_{3}=0\}$ and $\mathcal{F}\cap\{Z_{1}=1\}$, we have
(4.6) $\left.\left\langle\nabla\left(Z_{2}-Z_{3}\right),V\right\rangle\right|_{\{Z_{2}-Z_{3}=0\}}=4Z_{2}(X_{3}-X_{2})\geq-8Z_{2}(\sqrt{Z_{2}}-\sqrt{Z_{3}})=0,$
(4.7) $\begin{split}\left.\left\langle\nabla Z_{1},V\right\rangle\right|_{\{Z_{1}=1\}}&=2Z_{1}(X_{1}-X_{2})\leq 2Z_{1}\left(\sqrt{\frac{Z_{2}}{Z_{1}}}-\sqrt{Z_{1}Z_{2}}\right)=0.\end{split}$
Therefore, the vector field $V$ restricted on these faces points inward. For the face $\mathcal{F}\cap\left\{2(\sqrt{Z_{2}}-\sqrt{Z_{3}})+X_{3}-X_{2}=0\right\},$ we have
(4.8) $\left.\left\langle\nabla\left(2(\sqrt{Z_{2}}-\sqrt{Z_{3}})+X_{3}-X_{2}\right),V\right\rangle\right|_{\mathcal{F}\cap\left\{2(\sqrt{Z_{2}}-\sqrt{Z_{3}})+X_{3}-X_{2}=0\right\}}=(\sqrt{Z_{2}}-\sqrt{Z_{3}})K,$
where
$K:=1+(4m-4)\sqrt{Z_{3}}-4\sqrt{Z_{2}}+2Z_{1}(\sqrt{Z_{2}}+\sqrt{Z_{3}}).$
By [Chi26, Proposition 3.2], the factor $K$ is non-negative on $\mathcal{RS}\cap\{1-Z_{1}\geq 0\}\cap\{X_{3}\geq 0\}\cap\{Z_{2}-Z_{3}\geq 0\}$. Finally, by Proposition 4.1, we have $\left.\langle\nabla F_{2},V\rangle\right|_{\mathcal{F}\cap\{F_{2}=0\}}\geq 0$. Therefore, the set $\mathcal{F}$ is compact and invariant. ∎
In the following proposition, we show that the level sets $\{F_{l}=0\}$ serve as barriers within the 2-dimensional invariant sets of $\mathcal{RF}\cap\mathcal{RS}_{m=0}$ and $\mathcal{RF}\cap\mathcal{RS}_{\text{FS}}$, providing insight into the behavior of the integral curves that represent Ricci-flat metrics.a
Proposition 4.3.
The sets
(4.9) $\begin{split}\mathcal{U}_{+}:&=\mathcal{RF}\cap\mathcal{RS}_{m=0}\cap\{F_{2}\geq 0\}\cap\{X_{2}-\sqrt{Z_{1}Z_{2}}\geq 0\}\cap\{1-Z_{1}\geq 0\},\\ \mathcal{U}_{-}:&=\mathcal{RF}\cap\mathcal{RS}_{m=0}\cap\{F_{2}\leq 0\}\cap\{X_{2}-\sqrt{Z_{1}Z_{2}}\leq 0\},\\ \mathcal{D}_{+}:&=\mathcal{RF}\cap\mathcal{RS}_{\text{FS}}\cap\{F_{2m+2}\geq 0\}\cap\{X_{2}-\sqrt{Z_{1}Z_{2}}\geq 0\}\cap\{1-Z_{1}\geq 0\},\\ \mathcal{D}_{-}:&=\mathcal{RF}\cap\mathcal{RS}_{\text{FS}}\cap\{F_{2m+2}\leq 0\}\cap\{X_{2}-\sqrt{Z_{1}Z_{2}}\leq 0\}\end{split}$
are invariant. The integral curve $\xi(2,\pi,0,0)$ is the invariant set $\mathcal{U}_{0}:=\mathcal{U}_{+}\cap\mathcal{U}_{-}$. The integral curve $\xi(2m+2,0,0,0)$ is the invariant set $\mathcal{D}_{0}:=\mathcal{D}_{+}\cap\mathcal{D}_{-}$.
Proof.
Since $H=1$ and $X_{3}=0$ in $\mathcal{RF}\cap\mathcal{RS}_{m=0}$, the equality $X_{1}=1-2X_{2}$ holds. By (4.1), we have
(4.10) $\begin{split}\left.\langle\nabla F_{2},V\rangle\right|_{\mathcal{U}_{\pm}\cap\{F_{2}=0\}}&=\left(2\sqrt{\frac{Z_{2}}{Z_{1}}}(1-X_{1})-4Z_{2}\right)(1-Z_{1})\\ &=4(X_{2}-\sqrt{Z_{1}Z_{2}})\sqrt{\frac{Z_{2}}{Z_{1}}}(1-Z_{1}).\end{split}$
On $\mathcal{U}_{\pm}\cap\{X_{2}-\sqrt{Z_{1}Z_{2}}=0\}$, we have $F_{2}=3X_{2}-1+2\left(\sqrt{\frac{Z_{2}}{Z_{1}}}-\sqrt{Z_{1}Z_{2}}\right)=\sqrt{Z_{1}Z_{2}}-1+2\sqrt{\frac{Z_{2}}{Z_{1}}}$. Therefore,
(4.11) $\begin{split}\left.\langle\nabla(X_{2}-\sqrt{Z_{1}Z_{2}}),V\rangle\right|_{\mathcal{U}_{\pm}\cap\{X_{2}-\sqrt{Z_{1}Z_{2}}=0\}}&=-X_{2}+R_{2}-\sqrt{Z_{1}Z_{2}}(X_{1}-2X_{2})\\ &=-\sqrt{Z_{1}Z_{2}}+4Z_{2}-2Z_{1}Z_{2}-\sqrt{Z_{1}Z_{2}}(1-4\sqrt{Z_{1}Z_{2}})\\ &=2\sqrt{Z_{1}Z_{2}}F_{2}.\end{split}$
Since $F_{2}\geq 0$, the inequality (4.7) still holds in $\mathcal{U}_{+}$ by (4.10) and (4.11). Therefore, the set $\mathcal{U}_{+}$ is invariant. By (4.7) and (4.10), the set $\mathcal{U}_{0}$ is a 1-dimensional invariant set. By (3.12), the integral curve $\xi(2,\pi,0,0)$ is tangent to $\mathcal{U}_{0}$. Therefore, the integral curve is $\mathcal{U}_{0}$. By (4.11), it is clear that $\mathcal{U}_{-}\cap\{1-Z_{1}\leq 0\}$ is invariant. If an integral curve in $\mathcal{U}_{-}$ escapes, it has to transversally intersect $\mathcal{U}_{0}$, a contradiction. The set $\mathcal{U}_{-}$ is thus invariant.
The statements on $\mathcal{D}_{\pm}$ are proved with similar computations. Since $H=1$ and $X_{2}-X_{3}=0$ in $\mathcal{RF}\cap\mathcal{RS}_{\text{FS}}$, the equality $X_{1}=1-(4m+2)X_{2}$ holds. With $Z_{2}-Z_{3}=0$, we have analogous equations to (4.10) and (4.11). The integral curve $\xi(2m+2,0,0,0)$ is thus identified as an algebraic curve $\mathcal{D}_{0}$. ∎
Lemma 4.4.
Integral curves in $\left\{\xi(k,\theta,s_{4},s_{5})\mid k\in\{1,2\},\theta\in[0,\pi],s_{4},s_{5}\geq 0\right\}$ are defined on $\mathbb{R}$. Metrics represented by these integral curves with $\theta<\pi$ are complete.
Proof.
At $p_{0}$, the inequalities $1-Z_{1}>0$ and $X_{1}>0$ hold. The other defining inequalities in (4.5) vanish at $p_{0}$. By (3.12) and (3.16), the defining inequalities
$Z_{2}-Z_{3}\geq 0,\quad 2(\sqrt{Z_{2}}-\sqrt{Z_{3}})+X_{3}-X_{2}\geq 0,\quad X_{2}\geq 0,\quad X_{3}\geq 0$
hold near $p_{0}$.
We focus on the sign of $F_{2}$ near $p_{0}$. By the linearized solution, we have $\lim\limits_{\eta\to\infty}F_{2}=-1+\frac{2}{k}$. If $k=1$, the function is positive initially. Consider $k=2$. By (4.1), we have
(4.12) $\begin{split}F_{2}^{\prime}&=F_{2}\left(G-\frac{\epsilon}{2}W-1+4\sqrt{Z_{1}Z_{2}}\right)\\ &\quad+(1-X_{1}-2\sqrt{Z_{1}Z_{2}})2\sqrt{\frac{Z_{2}}{Z_{1}}}(1-Z_{1})+4mZ_{3}(1-Z_{1}).\end{split}$
By (3.12), we also have
(4.13) $\begin{split}&G-\frac{\epsilon}{2}W-1+4\sqrt{Z_{1}Z_{2}}\\ &\sim\left(-(16m^{2}+24m)s_{1}-(8m^{2}+16m)\sqrt{2}s_{3}-2s_{4}-(8m+6)\frac{\epsilon}{2}s_{5}\right)e^{2\eta}+O(e^{(2+\delta)\eta}),\\ &1-X_{1}-2\sqrt{Z_{1}Z_{2}}\\ &\sim\left((8m^{2}+12m)s_{1}+(4m^{2}+8m)\sqrt{2}s_{3}+s_{4}+(4m+2)\frac{\epsilon}{2}s_{5}\right)e^{2\eta}+O(e^{(2+\delta)\eta}).\end{split}$
If $F_{2}<0$ initially, the derivative $F_{2}^{\prime}$ is initially non-negative by (4.13), a contradiction. Note that leading terms in (4.13) vanish simultaneously only if $(k,\theta,s_{4},s_{5})=(2,\pi,0,0)$. By Proposition 4.3, the integral curve $\xi(2,\pi,0,0)$ still stays in $\mathcal{F}$. Therefore, each integral curve in $\left\{\xi(k,\theta,s_{4},s_{5})\mid k\in\{1,2\},\theta\in[0,\pi],s_{4},s_{5}\geq 0\right\}$ is in the set $\mathcal{F}$, with $\xi(2,\pi,0,0)$ being the only one that lies on the boundary $\partial\mathcal{F}$. By the escape lemma, each integral curve in the family is defined on $\mathbb{R}$. ∎
4.2. $k\geq 3$
Since $\lim\limits_{\eta\to-\infty}F_{2}=-1+\frac{2}{k}$, the function $F_{2}$ is initially negative along $\xi(k,\theta,s_{4},s_{5})$ if $k\geq 3$. The compact set $\mathcal{F}$ is no longer applicable. We proceed by applying an alternative method to analyze this case.
Define the following sets.
(4.14) $\begin{split}&\mathcal{A}:=\mathcal{RS}\cap\{1-Z_{1}\geq 0\}\cap\{X_{1}-X_{2}\leq 0\}\cap\{Z_{2}-Z_{3}\geq 0\}\\ &\quad\cap\left\{2(\sqrt{Z_{2}}-\sqrt{Z_{3}})+X_{3}-X_{2}\geq 0\right\}\cap\{X_{1},X_{2},X_{3}\geq 0\},\\ &\mathcal{B}:=\mathcal{RS}\cap\{1-Z_{1}\geq 0\}\cap\{X_{1}-X_{2}\geq 0\}\cap\{Z_{2}-Z_{3}\geq 0\}\\ &\quad\cap\left\{2(\sqrt{Z_{2}}-\sqrt{Z_{3}})+X_{3}-X_{2}\geq 0\right\}\cap\{X_{1},X_{2},X_{3}\geq 0\},\\ &\mathcal{C}:=\mathcal{RS}\cap\{1-Z_{1}\leq 0\}\cap\{X_{1}-X_{2}\geq 0\}.\end{split}$
We have the following lemma.
Lemma 4.5.
Sets $\mathcal{A}$ and $\mathcal{B}$ are compact. Sets $\mathcal{A}$ and $\mathcal{C}$ are invariant. Each $\xi(k,\theta,s_{4},s_{5})$ stays in $\mathcal{B}$, or eventually enters $\mathcal{C}$ through $\mathcal{B}\cap\{1-Z_{1}=0\}\cap\{X_{1}-X_{2}>0\}$, or eventually enters $\mathcal{A}$ through $\mathcal{B}\cap\{1-Z_{1}>0\}\cap\{X_{1}-X_{2}=0\}$.
Proof.
By the inequalities $1-Z_{1}\geq 0$ and $Z_{2}-Z_{3}\geq 0$, it is clear that $\mathcal{A}$ and $\mathcal{B}$ are compact. By [Chi26, Proposition 3.2], the set $\mathcal{A}$ is invariant. Since
(4.15) $\begin{bmatrix}Z_{1}\\ X_{1}-X_{2}\end{bmatrix}^{\prime}=\begin{bmatrix}2Z_{1}(X_{1}-X_{2})\\ (X_{1}-X_{2})\left(G-\frac{\epsilon}{2}W-1\right)+(4Z_{2}+4mZ_{3})(Z_{1}-1)\end{bmatrix}$,
the set $\mathcal{C}$ is invariant.
By (3.19), each $\xi(k,\theta,s_{4},s_{5})$ is initially in $\mathcal{B}$. The integral curve is in the boundary $\partial\mathcal{B}$ only if $\theta=0$. By (4.6)-(4.8) and the fact that $R_{i}\geq 0$ in $\mathcal{B}$, if the integral escapes $\mathcal{B}$, it does not leave the set through the following faces:
$\mathcal{B}\cap\{Z_{2}-Z_{3}=0\},\quad\mathcal{B}\cap\left\{2(\sqrt{Z_{2}}-\sqrt{Z_{3}})+X_{3}-X_{2}=0\right\},\quad\mathcal{B}\cap\{X_{i}=0\}.$
In other words, if a $\xi(k,\theta,s_{4},s_{5})$ leaves $\mathcal{B}$ at some $\eta_{*}$, either $1-Z_{1}$ or $X_{1}-X_{2}$ vanishes at $\eta_{*}$. These two functions cannot vanish simultaneously as the set $\mathcal{RS}_{\text{round}}$ is invariant while $p_{0}\notin\mathcal{RS}_{\text{round}}$. If the integral curve leaves $\mathcal{B}$ through a point with $1-Z_{1}=0$ and $X_{1}-X_{2}>0$, it enters the invariant set $\mathcal{C}$. If the integral curve leaves $\mathcal{B}$ through a point with $1-Z_{1}>0$ and $X_{1}-X_{2}=0$, then it enters the invariant set $\mathcal{A}$. The proof is complete. ∎
Note that $\mathcal{F}\subset\mathcal{B}$. While $\mathcal{B}$ being a larger set that include $\xi(k,\theta,s_{4},s_{5})$ with any $k\geq 1$, the trade-off is losing the invariance of the set. By Lemma 4.5, it is necessary to determine the parameters for which the integral curve enters $\mathcal{A}$ or stays in $\mathcal{B}$. In the Ricci-flat case with $\theta\in\{0,\pi\}$, this question is straightforward to resolve, as demonstrated in the following proposition.
Proposition 4.6.
The integral curve $\xi(k,\pi,0,0)$ eventually enters $\mathcal{A}$ if $k=1$, eventually enters $\mathcal{C}$ if $k\geq 3$. The integral curve $\xi(k,0,0,0)$ eventually enters $\mathcal{A}$ if $k\in[1,2m+1]$, eventually enters $\mathcal{C}$ if $k\geq 2m+3$.
Proof.
Let $X:=X_{1}-X_{2}$. The equalities $X_{1}=\frac{1+2X}{3}$ and $X_{2}=\frac{1-X}{3}$ hold on $\mathcal{E}\cap\mathcal{RS}_{m=0}$. Furthermore, the system (2.8) restricted on the $2$-dimensional invariant set $\mathcal{RF}\cap\mathcal{RS}_{m=0}$ is equivalent to
(4.16) $\begin{bmatrix}X\\ Z_{1}\end{bmatrix}^{\prime}=\begin{bmatrix}X(G-1)+4Z_{2}(Z_{1}-1)\\ 2Z_{1}X\end{bmatrix}$
where $Z_{2}=\frac{1-G}{8-2Z_{1}}\geq 0$. By Proposition 4.3, the integral curve $\xi(2,\pi,0,0)$ joins the critical point $(1,0)$ and $(0,1)$. As the critical point $(0,1)$ is a saddle in (4.16), each $\xi(k,\pi,0,0)$ does not converge to this critical point if $k\neq 2$.
The integral curve $\xi(1,\pi,0,0)$ is initially in $\mathcal{B}\cap\mathcal{U}_{+}$. If $\xi(1,\pi,0,0)$ escapes $\mathcal{B}$ through points in $\mathcal{B}\cap\{1-Z_{1}=0\}\cap\{X_{1}-X_{2}>0\}$, the inequality $F_{2}<0$ holds at the intersection point, meaning the integral curve escapes $\mathcal{U}_{+}$ first. This is a contradiction to Proposition 4.3. If $\xi(1,\pi,0,0)$ stays in $\mathcal{B}\cap\mathcal{RF}\cap\mathcal{RS}_{m=0}$, the monotonicity of $Z_{1}$ forces the integral curve to converge to $(0,1)$, a contradiction. Therefore, the integral curve $\xi(1,\pi,0,0)$ eventually enters $\mathcal{A}$.
On the other hand, the integral curve $\xi(k,\pi,0,0)$ is initially in $\mathcal{B}\cap\mathcal{U}_{-}$ if $k\geq 3$. If the integral curve escapes $\mathcal{B}$ through points in $\mathcal{B}\cap\{1-Z_{1}>0\}\cap\{X_{1}-X_{2}=0\}$, the inequality $F_{2}>0$ holds at the intersection point, meaning the integral curve escapes $\mathcal{U}_{-}$ first. This is a contradiction to Proposition 4.3. If $\xi(k,\pi,0,0)$ stays in $\mathcal{B}\cap\mathcal{RF}\cap\mathcal{RS}_{m=0}$, the integral curve converges to $(0,1)$, a contradiction. Therefore, the integral curve $\xi(k,\pi,0,0)$ eventually enters $\mathcal{C}$ if $k\geq 3$.
By Proposition 4.3, the $2$-dimensional set $\mathcal{B}\cap\mathcal{RF}\cap\mathcal{RS}_{\text{FS}}$ is divided into two components by the integral curve $\xi(2m+2,0,0,0)$. The statement for $\xi(k,0,0,0)$ is proved with the similar argument. ∎
Proposition 4.7.
The function $\frac{H-1}{\sqrt{-Q}}$ monotonically decreases along each $\xi(k,\theta,s_{4},s_{5})$ with $s_{4}>0$ as long as the integral curve is in $\mathcal{B}$.
Proof.
We first claim
(4.17) $Q+1-H+\frac{\epsilon}{2}W(1-H)<0$
along each $\xi(k,\theta,s_{4},s_{5})$ in $\mathcal{B}$ with $s_{4}>0$. By (3.18), we have $Q+1-H+\frac{\epsilon}{2}W(1-H)\sim-s_{4}e^{\eta}$. Hence, the function is initially negative if $s_{4}>0$. By (2.10), we have
[…]
For $\epsilon=0$, the last term is non-positive in $\mathcal{RS}$. For $\epsilon=1$, note that $R_{s}\geq 0$ in $\mathcal{B}$ and hence $(n-1)\frac{\epsilon}{2}W=Q+1-G-R_{s}\leq 1$. Therefore, the last term is still non-positive. The inequality $Q+1-H+\frac{\epsilon}{2}W(1-H)<0$ holds.
We have
[…]
The proof is complete. ∎
Proposition 4.8.
Along each $\xi(k,\theta,s_{4},s_{5})$ with $s_{4}>0$ while the integral curve is in $\mathcal{B}$, there exists an $\eta_{*}$ and a $C_{1}>0$ small enough such that $\sqrt{-Q}\geq C_{1}$ and $1-H\geq C_{1}^{2}$ for $\eta>\eta_{*}$.
Proof.
By Proposition 4.7, there exists an small enough $C_{1}>0$ and an $\eta_{*}$ such that $\frac{1-H}{\sqrt{-Q}}\geq C_{1}$ for $\eta>\eta_{*}$, where $C_{1}=\left(\frac{1-H}{\sqrt{-Q}}\right)(\eta_{*})>0$. By (4.17), we have
(4.20) $\begin{split}0&\geq-\frac{\epsilon}{2}W(1-H)>Q+1-H\geq\sqrt{-Q}(C_{1}-\sqrt{-Q})\end{split}$
for $\eta>\eta_{*}$ while the integral curve is in $\mathcal{B}$. Therefore, the inequalities $\sqrt{-Q}\geq C_{1}$ and $1-H\geq C_{1}^{2}$ hold. ∎
We proceed to prove the following lemma, which is a generalization to [App17, Lemma 5.4].
Lemma 4.9.
For each fixed $k\geq 3$ and $\theta\in\left[0,\pi\right]$, there exists an $\tilde{\alpha}_{k,\theta}\geq 0$ such that each $\xi(k,\theta,s_{4},s_{5})$ with $s_{5}\geq 0$ stays in $\mathcal{B}$ or eventually enters $\mathcal{A}$ if $s_{4}\geq\tilde{\alpha}_{k,\theta}$. Each $\xi(k,\theta,s_{4},s_{5})$ with $s_{4}\geq\tilde{\alpha}_{k,\theta}$ and $s_{5}\geq 0$ is thus defined on $\mathbb{R}$.
Proof.
Initially the integral curve $\xi(k,\theta,s_{4},s_{5})$ is in $\mathcal{B}$. For $s_{4}>0$, the function $Q$ is initially negative. Furthermore, as $R_{i}\geq 0$ in $\mathcal{B}$, we have $Q\in(0,1)$ along $\xi(k,\theta,s_{4},s_{5})$ in $\mathcal{B}$. We have
(4.21) $\begin{split}\left(\frac{\sqrt{Z_{1}}^{\prime}}{\sqrt{-Q}}\right)^{\prime}&=\left(\frac{\sqrt{Z_{1}}^{\prime}}{\sqrt{-Q}}\right)(X_{1}-X_{2}-1)+\frac{R_{1}-R_{2}}{\sqrt{-Q}}+\frac{\frac{\epsilon}{2}W(1-H)\sqrt{Z_{1}}^{\prime}}{Q\sqrt{-Q}}\\ &\leq\left(\frac{\sqrt{Z_{1}}^{\prime}}{\sqrt{-Q}}\right)(X_{1}-X_{2}-1)\\ &\leq\left(\frac{\sqrt{Z_{1}}^{\prime}}{\sqrt{-Q}}\right)(H-1)\leq 0.\end{split}$
The function $\frac{\sqrt{Z_{1}}^{\prime}}{\sqrt{-Q}}$ monotonically decreases while $\xi(k,\theta,s_{4},s_{5})$ is in $\mathcal{B}$.
By Proposition 4.8, there exists a small enough $\eta_{*}$ such that
(4.22) $\begin{split}\left(\frac{\sqrt{Z_{1}}^{\prime}}{\sqrt{-Q}}\right)^{\prime}&\leq-\left(\frac{\sqrt{Z_{1}}^{\prime}}{\sqrt{-Q}}\right)C_{1}^{2},\quad C_{1}=\left(\frac{1-H}{\sqrt{-Q}}\right)(\eta_{*})\end{split}$
holds along $\xi(k,\theta,s_{4},s_{5})$ in $\mathcal{B}$ for $\eta>\eta_{*}$. We have
(4.23) $\frac{\sqrt{Z_{1}}^{\prime}}{\sqrt{-Q}}\leq C_{2}e^{-C_{1}^{2}(\eta-\eta_{*})},\quad C_{2}=\left(\frac{\sqrt{Z_{1}}^{\prime}}{\sqrt{-Q}}\right)(\eta_{*})$
By (3.12) and the monotonicity of $\frac{\sqrt{Z_{1}}^{\prime}}{\sqrt{-Q}}$, we have
(4.24) $\begin{split}C_{2}&\leq\lim\limits_{\eta\to-\infty}\left(\frac{\sqrt{Z_{1}}^{\prime}}{\sqrt{-Q}}\right)(\xi(k,\theta,s_{4},s_{5}))\\ &=\lim\limits_{\eta\to-\infty}\left(\frac{\sqrt{Z_{1}}(X_{1}-X_{2})}{\sqrt{-Q}}\right)(\xi(k,\theta,s_{4},s_{5}))=\frac{(\sqrt{s_{1}}+\sqrt{s_{2}})k}{\sqrt{2s_{4}}}.\end{split}$
By (3.16), we have $C_{2}\leq\frac{k}{\sqrt{s_{4}}}$. It follows that
(4.25) $\begin{split}\sqrt{Z_{1}}^{\prime}&\leq\frac{k}{\sqrt{s_{4}}}\sqrt{-Q}e^{-C_{1}^{2}(\eta-\eta_{*})}\leq\frac{k}{\sqrt{s_{4}}}e^{-C_{1}^{2}(\eta-\eta_{*})}\end{split}$
Integrating (4.25) for both sides, we obtain
(4.26) $\begin{split}\sqrt{Z_{1}}&\leq\frac{k}{\sqrt{s_{4}}}\frac{1-e^{-C_{1}^{2}(\eta-\eta_{*})}}{C_{1}^{2}}+(\sqrt{Z_{1}})(\eta_{*})\leq\frac{k}{\sqrt{s_{4}}}\frac{1}{C_{1}^{2}}+(\sqrt{Z_{1}})(\eta_{*})\end{split}$
for $\eta>\eta_{*}$ while $\xi(k,\theta,s_{4},s_{5})$ is in $\mathcal{B}$. By (3.17), we have
$C_{1}^{2}=\left(\frac{(1-H)^{2}}{-Q}\right)(\eta_{*})=\frac{s_{4}}{2}e^{2\eta_{*}}+O(e^{(2+\delta)\eta_{*}}).$
Choose a small enough $\eta_{*}$ and a large enough $s_{4}$, the inequality $\sqrt{Z}_{1}\leq 1$ holds along $\xi(k,\theta,s_{4},s_{5})$ while the integral curve is in $\mathcal{B}$. Therefore, the integral curve stays in $\mathcal{B}$, or eventually enters $\mathcal{A}$ by Lemma 4.5. The existence of $\tilde{\alpha}_{k,\theta}$ is established. ∎
For each $k\geq 3$ and $\theta\in[0,\pi]$, define
(4.27) $\alpha_{k,\theta}$ $=\inf\left\{\tilde{\alpha}_{k,\theta}\mid\text{$\xi(k,\theta,s_{4},0)$ does not enter $\mathcal{C}$ for any $s_{4}\geq\tilde{\alpha}_{k,\theta}$}\right\},$
[…]
By Lemma 4.9, such $\alpha_{k,\theta}$ and $\beta_{k,\theta}$ exist. It is clear that $\beta_{k,\theta}\geq\alpha_{k,\theta}$.
Lemma 4.10.
For each fixed $k\geq 3$ and $\theta\in\left[0,\pi\right]$, the integral curve $\xi(k,\theta,\beta_{k,\theta},s_{5})$ does not enters $\mathcal{C}$ for any $s_{5}\geq 0$. The integral curve $\xi(k,\theta,\alpha_{k,\theta},0)$ stays in $\mathcal{B}$ if $\alpha_{k,\theta}>0$.
Proof.
We first prove the statement for $\beta_{k,\theta}$. Suppose otherwise, there exists ${s_{5}}_{*}\geq 0$ such that $\xi(k,\theta,\beta_{k,\theta},{s_{5}}_{*})$ eventually enters $\mathcal{C}$. By the the continuous dependence on parameters, there exists a small $\delta>0$ such that $\xi(k,\theta,\beta_{k,\theta}+\delta,s_{5*})$ eventually enters $\mathcal{C}$, a contradiction to the definition of $\beta_{k,\theta}$.
Assume $\alpha_{k,\theta}>0$. The argument for $\beta_{k,\theta}$ carries over so that $\xi(k,\theta,\alpha_{k,\theta},0)$ does not enters $\mathcal{C}$. If $\xi(k,\theta,\alpha_{k,\theta},0)$ enters $\mathcal{A}$, there exists a small $\delta>0$ such that $\alpha_{k,\theta}-\delta>0$ and $\xi(k,\theta,s_{4},0)$ does not enter $\mathcal{C}$ and enters $\mathcal{A}$ if $s_{4}\in(\alpha_{k,\theta}-\delta,\alpha_{k,\theta}]$, a contradiction to the definition of $\alpha_{k,\theta}$. ∎
Lemma 4.11.
If $k\geq 3$, there are infinitely many $\theta\in(0,\pi)$ such that $\xi(k,\theta,s_{4},0)$ stays in $\mathcal{B}$ for some $s_{4}>0$. If $k\in[3,2m+1]$, there exists some $\theta_{\star}\in(0,\pi)$ such that $\xi(k,\theta_{\star},0,0)$ stays in $\mathcal{B}$ and is defined on $\mathbb{R}$.
Proof.
Consider $k\geq 3$. Define $\alpha:=\max\limits_{\theta\in[0,\pi]}\{\alpha_{k,\theta}\}$. Each integral curve in the family
$\left\{\xi(k,\theta,s_{4},0)\mid\theta\in[0,\pi],s_{4}\geq\alpha\right\}$
enters $\mathcal{A}$ or stays in $\mathcal{B}$. By Proposition 4.6, the integral curve $\xi(k,\pi,0,0)$ enters $\mathcal{C}$, and thus $\alpha\geq\alpha_{k,0}>0$. Suppose $\xi(k,\theta,\alpha,0)$ enters $\mathcal{A}$. Consider the continuous 1-parameter family
$\left\{\xi\left(k,(1-\tau)\pi+\tau\theta,\tau\alpha,0\right)\mid\tau\in[0,1]\right\}.$
There exists some $\tau_{\theta}\in(0,1)$ such that $\xi_{\star}(k,\theta):=\xi\left(k,(1-\tau_{\theta})\pi+\tau_{\theta}\theta,\tau_{\theta}\alpha,0\right)$ stays in $\mathcal{B}$ by the continuous dependence on parameter. Since $\xi_{\star}(k,\theta_{1})\neq\xi_{\star}(k,\theta_{2})$ if $\theta_{1}\neq\theta_{2}$, there are infinitely many such $\xi_{\star}(k,\theta)$.
If $k\in[3,2m+1]$, the integral curve $\xi(k,0,0,0)$ enters $\mathcal{A}$, and $\xi(k,\pi,0,0)$ enters $\mathcal{C}$ by Proposition 4.6. Consider the continuous 1-parameter family $\left\{\xi\left(k,\theta,0,0\right)\mid\theta\in[0,\pi]\right\}.$ There exists some $\theta_{\star}\in(0,\pi)$ such that $\xi(k,\theta_{\star},0,0)$ stays in $\mathcal{B}$ by the continuous dependence on parameter. The integral curve is defined on $\mathbb{R}$ by the compactness of $\mathcal{B}$. ∎
5. Asymptotics
5.1. Non-Einstein Ricci solitons
If an integral curve enters $\mathcal{A}$, the analysis in [Chi26, Section 4] applies. We have the following lemma.
Lemma 5.1.
Each $\xi(k,\theta,s_{4},s_{5})$ with $s_{4},s_{5}>0$ is AC. Each $\xi(k,\theta,s_{4},0)$ with $s_{4}>0$ that enters $\mathcal{A}$ is ACP. The base of the asymptotic paraboloid is the non-Kähler $\mathbb{CP}^{2m+1}/\mathbb{Z}_{k}$ if $\theta\in(0,\pi)$.
We investigate the asymptotic for integral curves that stays in $\mathcal{B}$. The following proposition is extended directly from [Chi26, Section 4].
Proposition 5.2.
A $\xi(k,\theta,s_{4},s_{5})$ with $s_{4}>0$ and $s_{5}\geq 0$ converges to $(0,0,0,\mu^{2},0,0,0,0)$ for some $\mu\in[0,1]$. The function $Q$ thus converges to 1-1.
Proof.
The proof of [Chi26, Proposition 4.1] applies verbatim to our case. ∎
Lemma 5.3.
Each $\xi(k,\theta,s_{4},0)$ with $s_{4}>0$ that stays in $\mathcal{B}$ is AP. The base of the asymptotic paraboloid is the Jensen $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$ if $\theta\in(0,\pi)$.
Proof.
If a $\xi(k,\theta,s_{4},0)$ stays in $\mathcal{B}$, the function $Z_{1}$ monotonically increases. Therefore, the limit $\lim\limits_{\eta\to\infty}\sqrt{Z_{1}}=\mu^{2}$ exists for some $\mu\in(0,1]$ by Proposition 5.2. Apply [Win23, Proposition 2.22], the limit $\lim\limits_{\eta\to\infty}\sqrt{\frac{Z_{3}}{Z_{2}}}=\nu^{2}$ also exists. By [Chi26, Lemma 4.7], the possible values for $(\mu^{2},\nu^{2})$ are $(1,1)$ and $\left(1,\frac{1}{2m+3}\right)$, and we have the limits
(5.1) $\begin{split}\lim\limits_{t\to\infty}a\dot{a}&=\lim\limits_{\eta\to\infty}\frac{Q}{C}\frac{X_{1}Z_{1}}{Z_{2}}=\frac{1}{-C}\left(2+4m\nu^{4}\right),\\ \lim\limits_{t\to\infty}b\dot{b}&=\lim\limits_{\eta\to\infty}\frac{Q}{C}\frac{X_{2}}{Z_{2}}=\frac{1}{-C}\left(2+4m\nu^{4}\right),\\ \lim\limits_{t\to\infty}c\dot{c}&=\lim\limits_{\eta\to\infty}\frac{Q}{C}\frac{X_{3}}{Z_{4}}=\frac{1}{-C}\left((4m+8)-6\nu^{2}\right).\end{split}$
Therefore, the metric represented by $\xi(k,\theta,s_{4},0)$ is AP, with
(5.2) $\begin{split}a^{2}&\sim\frac{1}{\sqrt{-C}}\left(2+4m\nu^{4}\right)t,\\ b^{2}&\sim\frac{1}{\sqrt{-C}}\left(2+4m\nu^{4}\right)t,\\ c^{2}&\sim\frac{1}{\sqrt{-C}}\left((4m+8)-6\nu^{2}\right)t.\end{split}$
as $t\to\infty$. The paraboloid is the standard $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$ ($\nu^{2}=1$) or the Jensen $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$ ($\nu^{2}=\frac{1}{2m+3}$). If $\theta=0$, the integral curve $\xi(k,0,s_{4},0)$ stays in $\mathcal{RS}_{\text{FS}}$, where $X_{2}=X_{3}$ and $Z_{2}=Z_{3}$. The base of the asymptotic paraboloid is round. Since $\lim\limits_{\eta\to-\infty}\frac{Z_{3}}{Z_{2}}<1$ initially along $\xi(k,\theta,s_{4},0)$ with $\theta>0$, the limit $\nu^{2}\neq 1$ by [Chi26, Proposition 4.9]. The base of the asymptotic paraboloid is the Jensen $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$. ∎
5.2. Einstein metrics
We discuss asymptotics for integral curve that represent Ricci-flat metrics in this section. Recall that integral curves that represent Einstein metrics are on the invariant set $\mathcal{E}$, where the function H1H-1 and $Q$ identically vanish.
Lemma 5.4.
Each $\xi(k,\theta,0,s_{5})$ with $s_{5}>0$ that is defined on $\mathbb{R}$ is AH. Each $\xi(k,\theta,0,0)$ that enters $\mathcal{A}$ is ALC. The base of the asymptotic cone is the non-Kähler $\mathbb{CP}^{2m+1}/\mathbb{Z}_{k}$ if $\theta\in(0,\pi)$.
Proof.
By [Chi21, Section 5], it is known that each $\xi(k,\theta,0,0)$ converges to
$q_{1}=\left(0,\frac{1}{n-1},\frac{1}{n-1},0,\frac{1}{(n-1)^{2}}\frac{n-2}{n+1},\frac{1}{(n-1)^{2}}\frac{n-2}{n+1},\frac{1}{(n-1)^{2}}\frac{n-2}{n+1},0\right)\text{ if }\theta=0;$
$q_{2}=\left(0,\frac{1}{n-1},\frac{1}{n-1},0,(m+1)^{2}z_{*},z_{*},(m+1)z_{*},0\right),z_{*}=\frac{1}{(n-1)^{2}}\frac{n-2}{4(m^{2}+3m+1)}\text{ if }\theta\in(0,\pi);$
each $\xi(k,\theta,0,s_{5})$ with $s_{5}>0$ converges to the invariant set
$q_{0}=\left\{\left(\frac{1}{n},\frac{1}{n},\frac{1}{n},z_{1},0,0,0,\frac{2}{n\epsilon}\right)\mid z_{1}\in[0,1]\right\}.$
The AH asymptotic is established, since
$\lim\limits_{t\to\infty}\frac{\dot{a}}{a}=\lim\limits_{t\to\infty}\frac{\dot{b}}{b}=\lim\limits_{t\to\infty}\frac{\dot{c}}{c}=\sqrt{\frac{\epsilon}{2n}}.$
We focus on proving the asymptotic of $\xi(k,\theta,0,0)$ in $\mathcal{A}$. Since $\dot{b}=\frac{X_{2}}{\sqrt{Z_{2}}}$ and $\dot{c}=\frac{X_{3}}{\sqrt{Z_{4}}}$, the limits $b_{\infty}:=\lim\limits_{t\to\infty}\dot{b}$ and $c_{\infty}:=\lim\limits_{t\to\infty}\dot{c}$ exist. Furthermore, we have $\left(\frac{4}{b^{2}}+\frac{4mb^{2}}{c^{4}}\right)\sim\frac{n-2}{t^{2}}$ as $t\to\infty$.
To show $\lim\limits_{t\to\infty}a$ exists, consider
(5.3) $\begin{split}\ddot{\sqrt{Z_{1}}}=\dot{\sqrt{Z_{1}}}\left(-3\frac{\dot{b}}{b}-4m\frac{\dot{c}}{c}\right)-\sqrt{Z_{1}}(1-Z_{1})\left(\frac{4}{b^{2}}+\frac{4mb^{2}}{c^{4}}\right).\end{split}$
Since $\sqrt{Z_{1}}$ decreases to zero in $\mathcal{A}$, the inequality
(5.4) $\begin{split}\ddot{\sqrt{Z_{1}}}\leq-\dot{\sqrt{Z_{1}}}\frac{n(1+\delta)}{t}-\sqrt{Z_{1}}\frac{(n-2)(1-\delta)}{t^{2}}\end{split}$
eventually holds for some small $\delta>0$.
Let $c_{1}=n(1+\delta)$ and $c_{2}=(n-2)(1-\delta)$. Integrate both sides, the inequality
(5.5) $\begin{split}\frac{a}{b}=\sqrt{Z_{1}}&\leq K_{1}t^{\frac{-\left(c_{1}-1-\sqrt{(c_{1}-1)^{2}-4c_{2}}\right)}{2}}+K_{2}t^{\frac{-\left(c_{1}-1+\sqrt{(c_{1}-1)^{2}-4c_{2}}\right)}{2}}\end{split}$
eventually holds. Since the indexes respectively converge to 1-1 and (n2)-(n-2) as $\delta\to 0^{+}$, the inequality
(5.6) $\begin{split}\frac{a}{b}\leq K_{3}t^{-\frac{2}{3}}\end{split}$
eventually holds for some $K_{3}>0$.
Since
(5.7) $\ddot{a}=\left(-2\frac{\dot{b}}{b}-4m\frac{\dot{c}}{c}\right)\dot{a}+\left(\frac{a}{b}\right)^{3}\left(\frac{2}{b^{2}}+4m\frac{b^{2}}{c^{4}}\right)b,$
and $a$ is decreasing in $\mathcal{A}$, the inequality
(5.8) $\begin{split}\ddot{a}&\leq-\dot{a}\frac{(n-1)(1+\delta)}{t}+\left(\frac{a}{b}\right)^{3}\frac{(n-2)(1+\delta)b_{\infty}}{t}\\ &\leq-\dot{a}\frac{2(n-1)}{t}+K_{3}^{3}\frac{2(n-2)b_{\infty}}{t^{3}}\end{split}$
eventually holds. Multiply both sides by $t^{2(n-1)}$ and integrate, we have
[…]
eventually for some $K_{6}>0$. Therefore, the function $a$ is bounded above along an integral curve in $\mathcal{A}$. The ALC asymptotic is established. ∎
We investigate the asymptotic for a $\xi(k,\theta,0,0)$ that stays in $\mathcal{B}$ in the following.
Proposition 5.5.
If a $\xi(k,\theta,0,0)$ stays in $\mathcal{B}$, the integral curve converges to critical points $p_{1}$ or $p_{2}$, where
(5.10) $\begin{split}&p_{1}=\left(\frac{1}{n},\frac{1}{n},\frac{1}{n},1,\frac{1}{n^{2}},\frac{1}{n^{2}},\frac{1}{n^{2}},0\right),\quad p_{2}=\left(\frac{1}{n},\frac{1}{n},\frac{1}{n},1,(2m+3)^{2}z_{**},z_{**},(2m+3)z_{**},0\right),\\ &z_{**}=\frac{n-1}{n^{2}}\frac{1}{2(2m+3)^{2}+4m}.\end{split}$
Proof.
Since $X_{1}-X_{2}>0$ along an integral curve that stays in $\mathcal{B}$, the function $Z_{1}$ monotonically increases to some positive number. Consider the derivative
$\left(\frac{Z_{2}^{2m+3}Z_{3}^{2m}}{Z_{1}}\right)^{\prime}=2\frac{Z_{2}^{2m+3}Z_{3}^{2m}}{Z_{1}}(nG-H)=2\frac{Z_{2}^{2m+3}Z_{3}^{2m}}{Z_{1}}(nG-1)\geq 0.$
The function $\frac{Z_{2}^{2m+3}Z_{3}^{2m}}{Z_{1}}$ thus converges as $Z_{2}$ and $Z_{3}$ are bounded in $\mathcal{B}$ while $Z_{1}$ is bounded away from zero. As the $\omega$-limit set is invariant, the $\omega$-limit set of a $\xi(k,\theta,0,0)$ that stays in $\mathcal{B}$ is contained in $\mathcal{E}\cap\{nG=1\}.$
Again by the invariance of the $\omega$-limit set, the set must be contained in $\{p_{1},p_{2}\}$. We claim that $p_{1}$ and $p_{2}$ cannot both be in the $\omega$-limit set of the integral curve. Suppose otherwise, there exists a sequence $\{a_{k}\}_{k=1}^{\infty}$ with $\lim\limits_{k\to\infty}a_{k}=\infty$ such that $\left\{\left(\sqrt{\frac{Z_{3}}{Z_{2}}}\right)(a_{k})\right\}_{k=1}^{\infty}$ is bounded away from $1$ and $\frac{1}{2m+3}$. Then the $\omega$-limit set contains a point that is not in $\{p_{1},p_{2}\}$, a contradiction. ∎
Proposition 5.6.
Consider a $\xi(k,\theta,0,0)$ that stays in $\mathcal{B}$. Suppose the function $X_{2}-X_{3}$ is not identically zero, then $\xi(k,\theta,0,0)$ converges to $p_{2}$.
Proof.
By Proposition 5.5, it suffices to show that $\xi(k,\theta,0,0)$ does not converge to $p_{1}$. Suppose otherwise in the following.
If $X_{2}-X_{3}$ changes sign infinitely many times, it is necessary that $R_{2}-R_{3}$ changes sign infinitely many times. But the function $R_{2}-R_{3}$ is eventually negative if the integral curve converges to $p_{1}$, a contradiction.
Assume $X_{2}-X_{3}$ is eventually negative. By $\left(\frac{Z_{3}}{Z_{2}}\right)^{\prime}=4\frac{Z_{3}}{Z_{2}}(X_{2}-X_{3})$, the function $\frac{Z_{3}}{Z_{2}}$ is eventually decreasing. As $\frac{Z_{3}}{Z_{2}}=1$ at $p_{2}$ and $\frac{Z_{3}}{Z_{2}}\leq 1$ in $\mathcal{B}$, the function $\frac{Z_{3}}{Z_{2}}$ must converge to $\frac{1}{2m+3}$. Hence, the integral curve converges to $p_{2}$.
If $X_{2}-X_{3}$ is eventually positive, the function $\frac{Z_{3}}{Z_{2}}$ eventually monotonically increases. Suppose $\lim\limits_{\eta\to\infty}\frac{Z_{3}}{Z_{2}}=1$. We have
[…]
The coefficients for $X_{2}-X_{3}$ and $\sqrt{Z_{2}}-\sqrt{Z_{3}}$ above are positive at $p_{1}$. With the assumption that $X_{2}-X_{3}$ is eventually positive, the function $2(\sqrt{Z_{2}}-\sqrt{Z_{3}})+X_{3}-X_{2}$ eventually increases. But the function is non-negative in $\mathcal{B}$ and it vanishes at $p_{1}$. We reach a contradiction. ∎
Lemma 5.7.
A $\xi(k,\theta,0,0)$ that stays in $\mathcal{B}$ is AC. If $\theta=0$, then the base of the cone is the standard $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$. If $\theta\in(0,\pi)$, then the base of the cone is Jensen $\mathbb{S}^{4m+3}/\mathbb{Z}_{k}$.
Proof.
If a $\xi(k,\theta,0,0)$ stays in $\mathcal{B}$, the function $X_{1}-X_{2}$ stays positive along the integral curve. Then $Z_{1}$ monotonically increases to 11 from Proposition 5.5. If $\theta=0$, then the integral curve is in $\mathcal{E}\cap\mathcal{RS}_{\text{FS}}$ and it converges to $p_{1}$. If $\theta>0$, the integral curve converges to $p_{2}$ by Proposition 5.6. ∎
We are ready to prove the main theorems.
Proof for Theorem 1.1.
The completeness of Ricci solitons in Theorem 1.1 is proved by Lemma 4.4.
We show that ach $\xi(k,\theta,s_{4},0)$ in Lemma 4.4 enters $\mathcal{F}\cap\mathcal{A}$. Since $1-Z_{1}\geq 0$ holds in the invariant set $\mathcal{F}$, by Lemma 4.5, each $\xi(k,\theta,s_{4},0)$ in Lemma 4.4 enters $\mathcal{A}$ or stays in $\mathcal{B}$. We show that the latter situation is impossible.
Suppose an integral curve in Lemma 4.4 stays in $\mathcal{B}$. By (4.12), we have
(5.12) $\begin{split}F_{2}^{\prime}&=F_{2}G+(4\sqrt{Z_{1}Z_{2}}-1)F_{2}+(1-X_{1}-2\sqrt{Z_{1}Z_{2}})2\sqrt{\frac{Z_{2}}{Z_{1}}}(1-Z_{1})+4mZ_{3}(1-Z_{1})\\ &\geq(4\sqrt{Z_{1}Z_{2}}-1)F_{2}+(1-X_{1}-2\sqrt{Z_{1}Z_{2}})2\sqrt{\frac{Z_{2}}{Z_{1}}}(1-Z_{1})\\ &=(4\sqrt{Z_{1}Z_{2}}-1)\left(X_{2}-X_{1}\right)+\left(1-\frac{1}{2}\sqrt{\frac{X_{1}Z_{1}}{Z_{2}}}\frac{\sqrt{X_{1}}}{Z_{1}}\right)4Z_{2}(1-Z_{1})\end{split}$
If a $\xi(k,\theta,0,0)$ stays in $\mathcal{F}\cap\mathcal{B}$, the integral curve converges to $p_{1}$ or $p_{2}$ by Proposition 5.5. As the first term vanishes, while the factor $\left(1-\frac{1}{2}\sqrt{\frac{X_{1}Z_{1}}{Z_{2}}}\frac{\sqrt{X_{1}}}{Z_{1}}\right)>0$ at both critical points, the derivative $F_{2}^{\prime}$ is positive near $p_{1}$ and $p_{2}$ while the integral curve is in $\mathcal{A}$. But $F_{2}$ vanishes at $p_{1}$ and $p_{2}$. We reach a contradiction.
If a $\xi(k,\theta,s_{4},0)$ with $s_{4}>0$ stays in $\mathcal{F}\cap\mathcal{B}$, the first term is eventually positive by Proposition 5.2. By [Chi26, Lemma 4.7], the function $Z_{1}$ converges to 1 and the second term is eventually positive by Proposition 5.1. Since $F_{2}>0$ initially and the function vanishes at $(0,0,0,1,0,0,0,0)$, we reach a contradiction.
The asymptotics are thus established by Lemma 5.1 and Lemma 5.4. ∎
Proof for Theorem 1.2-1.4.
For Theorem 1.2-1.3, the completeness is proved by Lemma 4.9. If the integral curve enters $\mathcal{A}$, its asymptotic is either ACP by 5.1, or ALC by Lemma 5.4. If the integral curve enters $\mathcal{B}$, the Ricci soliton is non-collapsed. Its asymptotic is either AP by Lemma 5.3, or AC by Lemma 5.7.
To prove Theorem 1.4, by Lemma 5.3 and Lemma 5.7, it suffices to prove the existence of some $\xi(k,\theta,s_{4},0)$ and $\xi(k,\theta,0,0)$ that stay in $\mathcal{B}$. This is immediate by Lemma 4.11. ∎